OpenStudy (anonymous):

PLEASE HELP! The question asks to graph each line and to construct a perpendicular segment through the given point. It also asks to find the distance from the point to the line. The given equation and point is: 2x-3y=-9, (2,0). Please tell me all the steps to solve this and maybe the answer? I'm just a little confused on how to find the distance between two perpendicular lines.

OpenStudy (jdoe0001):

what's the slope of 2x-3y=-9 ?

OpenStudy (anonymous):

The slope would be 2/3x

OpenStudy (jdoe0001):

well... if you solve for "y", you'd end up with \(\bf y = \cfrac{2}{3}x+3\) so the slope will be just 2/3 so we want to find a perpendicular line that THAT ONE that passes through (2, 0) firstly so, a perpendicular to that will have a NEGATIVE RECIPROCAL slope what does that mean? RECIPROCAL of that 3/2 NEGATIVIZE IT -3/2 so we know now that the line passes through ( 2, 0) and has a slope of -3/2 so, you can find the equation for that line by using the point-slope form that is \(\large{ slope=m=-\cfrac{3}{2}\qquad \qquad \begin{array}{lllll} &x_1&y_1\\ &(2\quad ,&0) \end{array}\\ \quad \\ y-y_1=m(x-x_1)}\)

OpenStudy (anonymous):

Oh thank you! But from there on, how do you find the distance between those two perpendicular lines? I know how to find the distance if they were parallel cause they hve the same slope, but in this one they have opposite slopes... So how would I find it?

OpenStudy (jdoe0001):

hmm what did you get for the equation of the line perpendicular to that?

OpenStudy (anonymous):

I got y= -3/2x-3

OpenStudy (jdoe0001):

once you get that 2nd equation you'd end up with a SYSTEM OF EQUATIONS the solution to a system of equations is where the graph of both lines in the equation meet so you'd get a "x" and a "y" coordinate for that point, which will be the solution so you'd need the distance between ( 2 , 0 ) and the SOLUTION to the system of equations and for that you can use the distance formula, \(\bf \text{distance between 2 points}\\ d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\)

OpenStudy (jdoe0001):

ok... well... I got ....let's see \(\bf y-0=-\cfrac{3}{2}(x-2)\implies y = -\cfrac{3}{2}x+3\) recall that minus times minus = plus

OpenStudy (jdoe0001):

\(\bf y = \cfrac{2}{3}x+3\\ \quad \\ y = -\cfrac{3}{2}x+3\qquad \textit{notice the equivalence}\quad thus\\ \quad \\ \cfrac{2}{3}x+3=-\cfrac{3}{2}x+3\implies \cfrac{2}{3}x+\cfrac{3}{2}x=3-3\implies \cfrac{4x+9x}{6}=0\\ \quad \\ 13x=0\implies \color{red}{x =0}\\ \quad \\ \quad \\ \textit{let's use that in either equation, say the 1st one}\\ y = \cfrac{2}{3}0+3\implies \color{red}{y = 3}\)

OpenStudy (jdoe0001):

so, using the substitution method to solve the system of equations, that is, substituting "y" value of the 1st one, to the 2nd equation, we get x = 0 then using our x = 0 in the 1st one, we get y = 3 so the solution to the system of equations is ( 0, 3) that's where the lines meet

OpenStudy (anonymous):

Ok and then we use the distance formula. Thank you so much, you are literally a life saver!

OpenStudy (jdoe0001):

so the graph looks about say 

OpenStudy (jdoe0001):

yw